# Spatiotemporal patterns in the active cyclic Potts model

### v1

@@ IV Dynamics for Asymmetric Flipping Energies
In this section, we describe the dynamics with asymmetric flipping energies, for symmetric contact energies $J_{\alpha\alpha}=2$. By asymmetric flipping, we mean that $h_{01}$, $h_{12}$, and $h_{20}$ are not all equal. We still assume, however, that $h_{01}+h_{12}+h_{20}=h_{\mathrm{cyc}}\neq 0$ and keep $h_{\beta\alpha}=-h_{\alpha\beta}$. First, we consider a small system of $N=128^{2}$ sites, in which the HC and SW modes can coexist in the symmetric condition. When the flipping energies slightly deviate from the symmetric condition, the dynamics exhibit only small changes. For example, the SW+HC mode is obtained at $h_{01}=1.1$, $h_{12}=h_{20}=1$, similar to that at $h_{01}=h_{12}=h_{20}=1$, and the mean densities of $s=0$ and $s=2$ are slightly lower and higher $1/3$ in the SW mode, respectively (see Fig. 4 and Movies S1 and S2).
@@
However, large asymmetry alters the dynamics qualitatively. In symmetric conditions, the three-phase contact points move isotropically, and the boundaries between domains move all at the same speed in the direction from $s=[\alpha+1]$ to $s=\alpha$. In the case of asymmetric flipping energies, the speeds of the different boundaries are not the same. In the SW mode, when the flipping energies differ substantially, the three-phase contact points move ballistically rather than diffusively (see Fig. 5 and Movie S3). For $h_{01}=1.6$, $h_{12}=1.7$, $h_{20}=1$, as in Fig. 5, small biphasic domains of $s=1$ and $2$ move like amoeba locomotion and bacteria colony growth in the direction of the $s=1$ region. Their large fluctuations often result in domain division and disappearance.
@@
In the asymmetric case, the average fraction $\rho_{\alpha}=N_{\alpha}/N$ of sites in each state differs from $1/3$, as can be seen in Fig. 5. The contact energies turn out to be essential in determining those fractions. Indeed, for $h_{01}=1.6$, $h_{12}=1.7$, and $h_{20}=1$, the mean-field approach developed in Sec. VI.1, which neglects both the contact energies and the spatial structures, predicts $\rho_{0}\simeq\rho_{1}\simeq 0.32$ and $\rho_{2}\simeq 0.37$, whereas for $J=2$ the actual ratio $\rho_{0}/\rho_{1}\approx 5$ (see Fig. 5). This is expected, since a single site flip within a domain involves the loss of four $J_{\alpha\alpha}=J$ contacts, whereas a boundary flip involves the loss of two contacts ($42J=$ 4). The nucleation of a domain and the motion of an interface are therefore cooperative events, and they strongly affect the state ratios.
@@
Dynamic phase diagrams are shown in Fig. 6. In addition to the SW and HC modes, new one-phase modes Eα, in which the state $s=\alpha$ is predominant as in equilibrium, appear when either $h_{01}$, $h_{12}$ or $h_{20}$ dominates. Once a one-phase mode Eα is installed, it does not transform into the SW mode nor into the next state $s=[\alpha+1]$ of the cycling process. Temporal coexistence SW+Eα are possible, as shown in Fig. 6. The various modes are distinguished as follows. When $p_{\mathrm{three}}>0.05$ (see its definition in Sec. III), the system is in the SW mode. When $p_{\mathrm{one}}>0.05$, the system is either in the HC mode or in one of the Eα modes. In the case both of them exceed $0.05$, the system is either in the coexistence mode SW+HC or in a coexistence mode SW+Eα. The HC and Eα modes are distinguished from the distribution of $N_{s}$ shown in Fig. 7. When a peak exists at $N_{\alpha}/N\approx 1$ and it is ten times higher than their local minima close to $N_{\alpha}/N=1$, we consider that the dominant phase of $s=\alpha$ exists in a recognizable period. When all three states take the dominant phase, it is in the HC mode. Otherwise, we consider it to be in the Eα mode, in which the $s=\alpha$ state has the maximum peak at $N_{s}/N\approx 1$. The distribution in Fig. 7(b) belongs to E2 but is close to the threshold of the HC mode.
@@
As $h_{01}$ increases while $h_{12}$ and $h_{20}$ are fixed, the dynamics changes from E0, to SW+E0, SW, SW+E2, and E2 (or HC instead of E0), in order (see Figs. 6 and 8). Interestingly, high $h_{01}$ generates the E2 mode, although $h_{01}$ directly enhances the nucleation and growth of the $s=1$ phase. This dynamics can be understood by seeing the lifetimes of one-state dominant phases shown in Fig. 8(c). The lifetime of $s=\alpha$ is calculated as the period during which the system stays in $N_{\alpha}>0.98N$. The mean lifetime of the $s=0$ dominant phase exponentially decreases with increasing $h_{01}$, by more frequent nucleation and growth of the $s=1$ phase in the $s=0$ phase, as expected. Conversely, the mean lifetime of the $s=1$ phase is almost independent of $h_{01}$, because the $s=1$ phase is terminated by the nucleation and growth of the $s=2$ phase, which is determined by $h_{12}$. However, the $s=2$ phase remains for longer periods with increasing $h_{01}$, owing to the suppression of the nucleation. To understand this, let us consider an isolated dimer of $s=0$ sites in the $s=2$ phase. One site in this dimer can go back to the $s=2$ state via two pathways. One is the direct flipping to $s=2$, whose probability is min$(1,\exp(2J-h_{20}))$, since the number of contacts between $s=0$ and $s=2$ sites changes from six to four. The other is the two-step flipping via $s=1$, in which the probabilities of $s=0\to 1$ and $1\to 2$ are min$(1,\exp(-J+h_{01}))$ and min$(1,\exp(3J+h_{12}))$, respectively. The latter flipping increases for higher $h_{01}$, since the first step is rate-limiting. Thus, the $s=2$ phase becomes temporally dominant. Conversely, the acceleration of the second step does not enhance the latter flipping, as seen in the constant lifetime of the $s=1$ state.
@@
A similar tendency in the flipping energies is also found in the SW mode. When $h_{\alpha[\alpha+1]}$ is the highest among the three flipping energies, the $s=[\alpha+2]$ state is the most occupied over the lattice (see Figs. 4 and 5). The lifetime of the SW mode is calculated as the lapse from the time entering $0.1N<N_{\alpha}<0.75N$ for all $\alpha$ to that entering the one phase mode ($N_{\alpha}>0.98N$ for one of the states). When the three forward flips are comparable (e.g., $h_{01}\simeq 1$ while $h_{12}=h_{20}=1$ as in Fig. 8), the mean lifetime of the SW mode becomes longer and the system falls into this mode. As the flipping energies deviate more from the symmetric condition, one of the states ($s=\alpha$) dominates the SW mode, and subsequently the SW mode changes into the $s=\alpha$ dominant phase, Eα, via the SW+ Eα mode or the HC mode (see Fig. 6).
@@
Figure 11: Effects of the asymmetric contact energies with the symmetric flipping energies $h=h_{01}=h_{12}=h_{20}$ at $J_{11}=J_{22}=2$ and $N=128^{2}$. (a)–(c) Dependence on the contact energy $J_{00}$ at $h=h_{01}=h_{12}=h_{20}=1.3$. The dynamic modes are described at the top of the figure. (a) Probabilities of three states. The gray up-pointing triangles or squares, blue down-pointing triangles or crosses, and red circles or diamonds represent the data for $s=0$, $1$, and $2$, respectively. (b) Probabilities of one-phase state and three-phase coexistence. (c) Lifetimes of the homogeneous phases ($s=0$, $1$, and $2$) and SW mode. The dashed lines in (a–c) represent kinetically trapped modes. (d) Dynamic phase diagram as a function of $J_{00}$. The squares (light blue) and circles (gray) represent the HC and SW modes, respectively. The up-pointing and down-pointing triangles (green) represent the homogeneous phases of $s=0$ and $s=2$ (E0 and E2), respectively. The pluses (blue), open diamonds (magenta), closed diamonds (magenta), and crosses (red) represent the temporal coexistence: SW+HC, SW+E0, SW+E1, SW+E2, respectively. The asterisks indicate the hysteresis of two modes (E0 and SW+E1 or E1).
@@ V Dynamics for Asymmetric Contact Energies
Figure 12: Coexistence of the SW and E0 modes at $J_{00}=2.5$, $J_{11}=J_{22}=2$, $h=h_{01}=h_{12}=h_{20}=1.3$, and $N=128^{2}$. (a) Sequential snapshots in the SW mode. A biphasic domain moves in the direction of the $s=2$ region. (b) Time evolution of numbers of three states (corresponding to snapshots in (a)). The light gray, blue (dark gray), and red (medium gray) colors represent the data for $s=0$, $1$, and $2$, respectively. (c) Time evolution of numbers of three states starting from a $s=0$ dominant phase (E0 mode).
Interestingly, hysteresis occurs at high $h$ and high $J_{00}$. Two different modes appear depending on the initial condition, as shown in Fig. 12 and Movie S4: the E0 mode and the SW+E1 mode. They do not change into each other in our simulation times $\sim 10^{8}$, since their lifetimes become longer due to high energy barriers for the transitions. In the $s=1$ phase, biphasic domains of $s=2$ and $s=0$ often appear but do not grow to cover the entire lattice (see Figs. 12(a) and (b)). At relatively low $h$, the SW+E1 branch discretely appears in the middle of the E0 region (see Fig. 11). Thus, the SW+E1 mode is likely a kinetically trapped metastable state but the E0 mode remains as the only solution in the long time limit. In contrast, at high $h$, the E0 branch is only connected with the E0-mode solution at high $J_{00}$, and the SW+E1 branch is continuously connected with the SW+E1-mode solution at low $J_{00}$ (see the region of $h\geq 1.5$ in Fig. 11(d)). The SW+E1 mode becomes the E1 mode at high $J_{00}$ ($J_{00}\geq 2.5$ at $h=1.5$). Thus, the stable mode in the long time limit likely changes from the SW+E1 or E1 mode to E0 mode in the middle of the hysteresis region.
@@
VI Theoretical Analysis
@@ VI.1.1 Equilibrium State and Free Energy
Figure 14: Continuous model with $h_{01}=h_{12}=h_{20}=0.3$, obtained from $\varepsilon_{0}=0$, $\varepsilon_{1}=-0.3$, $\varepsilon_{2}=-0.6$, and $h_{20}^{\mathrm{neq}}=h_{\mathrm{cyc}}=0.9$. The interaction and gradient coefficients are $u=4$ and $k=1$, respectively. (a) Free energy landscape $f$ as a function of $\rho_{1}$ and $\rho_{2}$. The bottom left, bottom right and top left minima correspond to the phases $s=0$, $s=1$, and $s=2$, respectively. (b) Numerical integration of Eqs. (12)–(14) in one dimension, showing a biphasic $s=2,1$ (red, blue) solitary wave in an $s=0$ (gray) background. Time flows from top to bottom. The time lapse between two snapshots: $4\times 10^{5}$ time steps $dt=1.25\times 10^{-4}$. The RGB color encodes the system state as $\rho_{0}\mathcal{G}+\rho_{1}\mathcal{B}+\rho_{2}\mathcal{R}$, with $\mathcal{G}$ gray, $\mathcal{B}$ blue, and $\mathcal{R}$ red. (c) Two different-sized bands moving in opposite directions, colliding and annihilating each other. The same parameters as in the previous image.
@@ VI.2 Continuum Theory with Spatial Gradients and Interactions
$\Delta H_{01}(\mathbf{x})$ $=\left.\frac{\delta F}{\delta\rho_{1}(\mathbf{x})}\right|_{\rho_{2}}=\varepsilon_{1}-\varepsilon_{0}-u_{1}\rho_{1}+u_{0}\rho_{0}$ (18)
$-k\nabla^{2}\left(\rho_{1}-\rho_{0}\right)-h_{01}^{\mathrm{neq}}=-\Delta H_{10}(\mathbf{x}).$
@@
Figure 15: Evolution of a biphasic $s=1,2$ band in a $s=0$ backgroud, for asymmetric flipping energies. (a) Weak asymmetry: $h_{01}=0.34$, $h_{12}=0.3$, and $h_{20}=0.3$, obtained from $\varepsilon_{0}=0$, $\varepsilon_{1}=-0.34$, $\varepsilon_{2}=-0.64$, and $h_{20}^{\mathrm{neq}}=h_{\mathrm{cyc}}=0.94$. (b) Large asymmetry: $h_{01}=0.4$, $h_{12}=0.45$, and $h_{20}=0.3$, obtained from $\varepsilon_{0}=0$, $\varepsilon_{1}=-0.4$, $\varepsilon_{2}=-0.85$, and $h_{20}^{\mathrm{neq}}=h_{\mathrm{cyc}}=1.15$.

### v2

@@ IV Dynamics for Asymmetric Flipping Energies
In this section, we describe the dynamics with asymmetric flipping energies, for symmetric contact energies $J_{\alpha\alpha}=2$. By asymmetric flipping, we mean that $h_{01}$, $h_{12}$, and $h_{20}$ are not all equal. We still assume, however, that $h_{01}+h_{12}+h_{20}=h_{\mathrm{cyc}}\neq 0$ and keep $h_{\beta\alpha}=-h_{\alpha\beta}$. First, we consider a small system of $N=128^{2}$ sites, in which the HC and SW modes can coexist in the symmetric condition. When the flipping energies slightly deviate from the symmetric condition, the dynamics exhibit only small changes. For example, the SW+HC mode is obtained at $h_{01}=1.1$, $h_{12}=h_{20}=1$, similar to that at $h_{01}=h_{12}=h_{20}=1$, and the mean densities of $s=0$ and $s=2$ are slightly lower and higher than $1/3$ in the SW mode, respectively (see Fig. 5 and Movies S1 and S2).
@@
However, large asymmetry alters the dynamics qualitatively. In symmetric conditions, the three-phase contact points move isotropically, and the boundaries between domains move all at the same speed in the direction from $s=[\alpha+1]$ to $s=\alpha$. In the case of asymmetric flipping energies, the speeds of the different boundaries are not the same. In the SW mode, when the flipping energies differ substantially, the three-phase contact points move ballistically rather than diffusively (see Fig. 6 and Movie S3). For $h_{01}=1.6$, $h_{12}=1.7$, and $h_{20}=1$, as in Fig. 6, small biphasic domains of $s=1$ and $2$ move like amoeba locomotion and bacteria colony growth in the direction of the $s=1$ region. Their large fluctuations often result in domain division and disappearance.
@@
In the asymmetric case, the average fraction $\rho_{\alpha}=N_{\alpha}/N$ of sites in each state differs from $1/3$, as can be seen in Fig. 6. The contact energies turn out to be essential in determining those fractions. Indeed, for $h_{01}=1.6$, $h_{12}=1.7$, and $h_{20}=1$, the mean-field approach developed in Sec. VI.1, which neglects both the contact energies and the spatial structures, predicts $\rho_{0}\simeq\rho_{1}\simeq 0.32$ and $\rho_{2}\simeq 0.37$, whereas for $J=2$ the actual ratio $\rho_{0}/\rho_{1}\approx 5$ (see Fig. 6). This is expected, since a single site flip within a domain involves the loss of four $J_{\alpha\alpha}=J$ contacts, whereas a boundary flip involves the loss of two contacts ($42J=$ 4). The nucleation of a domain and the motion of an interface are, therefore, cooperative events, and they strongly affect the state ratios.
@@
Dynamic phase diagrams are shown in Fig. 7. In addition to the SW and HC modes, new one-phase modes Eα, in which the state $s=\alpha$ is predominant as in equilibrium, appear when either $h_{01}$, $h_{12}$ or $h_{20}$ dominates. Temporal coexistence SW+Eα is possible, as shown in Fig. 7. The various modes are distinguished as follows. When $p_{\mathrm{three}}>0.05$ (see its definition in Sec. III), the system is in the SW mode. When $p_{\mathrm{one}}>0.05$, the system is either in the HC mode or in one of the Eα modes. In the case both of them exceed $0.05$, the system is either in the coexistence mode SW+HC or in a coexistence mode SW+Eα. The HC and Eα modes are distinguished from the distribution of $N_{s}$ shown in Fig. 8. When all three states have peaks at $N_{s}/N\approx 1$, the system is in the HC mode. Otherwise, we consider it to be in the Eα mode, in which the $s=\alpha$ state has the maximum peak at $N_{s}/N\approx 1$. Since a very small peak can be buried or formed by statistical error, we consider that a peak is recognizably high as a dominant phase, when the peak at $N_{\alpha}/N\approx 1$ is ten times higher than the local minimum close to $N_{\alpha}/N=1$ (see the lower peak of $s=0$ in Fig. 8(b), which is in the E2 mode near the threshold to the HC mode). Thus, the cycling of two homogeneous phases still continues near the mode boundary in the Eα region of the phase diagram, but the cycling completely disappears deeply in the Eα region (compare Fig. 8(a) and (b)).
@@
As $h_{01}$ increases while $h_{12}$ and $h_{20}$ are fixed, the dynamics changes from E0, to SW+E0, SW, SW+E2, and E2 (or HC instead of E0), in order (see Figs. 7 and 9). Interestingly, high $h_{01}$ generates the E2 mode, although $h_{01}$ directly enhances the nucleation and growth of the $s=1$ phase. This dynamics can be understood by seeing the lifetimes of one-state dominant phases shown in Fig. 9(c). The lifetime of $s=\alpha$ is calculated as the period during which the system stays in $N_{\alpha}>0.98N$. The mean lifetime of the $s=0$ dominant phase exponentially decreases with increasing $h_{01}$, by more frequent nucleation and growth of the $s=1$ phase in the $s=0$ phase, as expected. Conversely, the mean lifetime of the $s=1$ phase is almost independent of $h_{01}$, because the $s=1$ phase is terminated by the nucleation and growth of the $s=2$ phase, which is determined by $h_{12}$. However, the $s=2$ phase remains for longer periods with increasing $h_{01}$, owing to the suppression of the nucleation. To understand this, let us consider an isolated dimer of $s=0$ sites in the $s=2$ phase. One site in this dimer can go back to the $s=2$ state via two pathways. One is the direct flipping to $s=2$, whose probability is min$(1,\exp(2J-h_{20}))$, since the number of contacts between $s=0$ and $s=2$ sites changes from six to four. The other is the two-step flipping via $s=1$, in which the probabilities of $s=0\to 1$ and $1\to 2$ are min$(1,\exp(-J+h_{01}))$ and min$(1,\exp(3J+h_{12}))$, respectively. The latter flipping increases for higher $h_{01}$, since the first step is rate-limiting. Thus, the $s=2$ phase becomes temporally dominant. Conversely, the acceleration of the second step does not enhance the latter flipping, as seen in the constant lifetime of the $s=1$ state.
@@
A similar tendency in the flipping energies is also found in the SW mode. When $h_{\alpha[\alpha+1]}$ is the highest among the three flipping energies, the $s=[\alpha+2]$ state is the most occupied over the lattice (see Figs. 5 and 6). The lifetime of the SW mode is calculated as the lapse from the time entering $0.1N<N_{\alpha}<0.75N$ for all $\alpha$ to that entering the one phase mode ($N_{\alpha}>0.98N$ for one of the states). When the three forward flips are comparable (e.g., $h_{01}\simeq 1$ while $h_{12}=h_{20}=1$ as in Fig. 9), the mean lifetime of the SW mode becomes longer and the system falls into this mode. As the flipping energies deviate more from the symmetric condition, one of the states ($s=\alpha$) dominates the SW mode, and subsequently the SW mode changes into the $s=\alpha$ dominant phase, Eα, via the SW+ Eα mode or the HC mode (see Fig. 7).
@@
Figure 12: Effects of the asymmetric contact energies with the symmetric flipping energies $h=h_{01}=h_{12}=h_{20}$ at $J_{11}=J_{22}=2$ and $N=128^{2}$. (a)–(c) Dependence on the contact energy $J_{00}$ at $h=h_{01}=h_{12}=h_{20}=1.3$. The dynamic modes are described at the top of the figure. (a) Probabilities of three states. The gray up-pointing triangles or squares, blue down-pointing triangles or crosses, and red circles or diamonds represent the data for $s=0$, $1$, and $2$, respectively. (b) Probabilities of one-phase state and three-phase coexistence. The dashed lines in (a–b) represent kinetically trapped modes (SW+E1 or E1). (c) Lifetimes of the homogeneous phases ($s=0$, $1$, and $2$) and SW mode. (d) Dynamic phase diagram as a function of $J_{00}$ for various $h$. The squares (light blue) and circles (gray) represent the HC and SW modes, respectively. The up-pointing and down-pointing triangles (green) represent the homogeneous phases of $s=0$ and $s=2$ (E0 and E2), respectively. The pluses (blue), open diamonds (magenta), closed diamonds (magenta), and crosses (red) represent the temporal coexistence: SW+HC, SW+E0, SW+E1, SW+E2, respectively. The asterisks indicate the hysteresis of two modes (E0 and SW+E1 or E1).
@@ V Dynamics for Asymmetric Contact Energies
Figure 13: E1 and E0 modes in the hysteresis region at $J_{00}=2.5$, $J_{11}=J_{22}=2$, $h=h_{01}=h_{12}=h_{20}=1.3$, and $N=128^{2}$. (a) Sequential snapshots in the E1 mode. A biphasic domain moves in the direction of the $s=2$ region. The green lines in the last snapshot represent the trajectories of two biphasic domains consisting of $s=0$ and $2$ states. The right trajectory corresponds to the sequential snapshots in (a). The left trajectory corresponds to the domain shown in the last period of Movie S4. (b) Time evolution of the fraction of each state (corresponding to snapshots in (a)). The light gray, blue (dark gray), and red (medium gray) colors represent the data for $s=0$, $1$, and $2$, respectively. (c) Time evolution of the fraction of each state starting from a $s=0$ dominant phase (E0 mode).
Interestingly, hysteresis occurs at high $h$ and high $J_{00}$. In the corresponding region of Fig. 12(d) (see also Movie S4), either the E0 mode or the E1 (or SW+E1) mode appears depending on the initial conditions. Those modes do not change into each other in our simulation times $\sim 10^{8}$, since their lifetimes become longer due to high transition energy barriers. In the E1 mode, biphasic domains of $s=2$ and $s=0$ often appear but do not grow to cover the entire lattice (see Figs. 13(a) and (b)). Interestingly, for $1.1\leq h\leq 1.3$, the SW+E1 or E1 modes, although they can be stabilized inside the hysteresis region, disappear in favor of the $E_{0}$ mode on both sides of the hysteresis region (see Fig. 12(b)). Thus, the SW+E1 mode is likely a kinetically trapped metastable state, while the E0 mode seems to be the stable state. In contrast, at $h\geq 1.5$, the E0 mode continues outside of the hysteresis region at higher $J_{00}$, while it is the SW+E1 mode that continues on the other side at lower $J_{00}$, see Fig. 12(d). The SW+E1 mode becomes the E1 mode in the middle of the hysteresis region ($J_{00}\geq 2.5$ at $h=1.5$). Thus, the stable mode in the long-time limit likely changes from the SW+E1 or E1 mode to the E0 mode in the middle of the hysteresis region.
@@
The ballistic motion of biphasic domains is clearly seen in Fig. 13(a) and Movie S4. The trajectories of two domains are shown in the last snapshot of Fig. 13(a) (the right one corresponds to the sequential snapshots in Fig. 13(a), and the left top one to the last period of Movie S4). The right trajectory consists of two straight (ballistic) parts and a dangling (fluctuating at a location) part. The first straight part corresponds to the motion of the small biphasic domain shown in the first snapshot of Fig. 13(a), and the second corresponds to that of the large biphasic domain shown in the three middle snapshots of Fig. 13(a). Therefore, a domain moves ballistically when it is biphasic, but it stops when it becomes a single phase.
@@
VI Phenomenological mean-field model
@@ VI.2 Continuum Theory with Spatial Gradients and Interactions
Figure 15: Dynamics in the continuous model. (a)-(b) Symmetric case $h_{01}=h_{12}=h_{20}=0.3$, obtained from $\varepsilon_{0}=0$, $\varepsilon_{1}=-0.3$, $\varepsilon_{2}=-0.6$, and $h_{20}^{\mathrm{neq}}=h_{\mathrm{cyc}}=0.9$. The interaction and gradient coefficients are $u=4$ and $k=1$, respectively. (a) Free energy landscape $f$ as a function of $\rho_{1}$ and $\rho_{2}$. The bottom left, bottom right and top left minima correspond to the phases $s=0$, $s=1$, and $s=2$, respectively. (b) Numerical integration of Eqs. (12)–(14) in one dimension, showing two different-sized $s=2,1$ (red, blue) biphasic solitary waves moving in opposite directions, colliding and annihilating each other. Time flows from top to bottom. The time lapse between two snapshots: $4\times 10^{5}$ time steps $dt=1.25\times 10^{-4}$. The RGB color encodes the system state as $\rho_{0}\mathcal{G}+\rho_{1}\mathcal{B}+\rho_{2}\mathcal{R}$, with $\mathcal{G}$ gray, $\mathcal{B}$ blue, and $\mathcal{R}$ red. (c) Asymmetric case. Evolution of a biphasic $s=1,2$ band in a $s=0$ backgroud for asymmetric flipping energies: $h_{01}=0.4$, $h_{12}=0.45$, and $h_{20}=0.3$, obtained from $\varepsilon_{0}=0$, $\varepsilon_{1}=-0.4$, $\varepsilon_{2}=-0.85$, and $h_{20}^{\mathrm{neq}}=h_{\mathrm{cyc}}=1.15$. As the band propagates, the blue phase ($s=1$) recedes, and after its disappearance, the red phase ($s=2$) is in turn consumed by the grey phase ($s=0$).
@@
$\Delta H_{01}(\mathbf{x})$ $=$ $\left.\frac{\delta F}{\delta\rho_{1}(\mathbf{x})}\right|_{\rho_{2}}$ (18)
$=$ $\varepsilon_{1}-\varepsilon_{0}-u_{1}\rho_{1}+u_{0}\rho_{0}-k\nabla^{2}\left(\rho_{1}-\rho_{0}\right)-h_{01}^{\mathrm{neq}}$
$=$ $-\Delta H_{10}(\mathbf{x}).$